OpenStudy (anonymous):

Use an identity to write the expression as a single trig value tan34 degrees/ 2(1-tan^2 34degrees)

OpenStudy (jdoe0001):

http://www.freemathhelp.com/images/doubleangles.png <-- notice the tangent double-angle identity

OpenStudy (jdoe0001):

\(\bf \cfrac{2tan(\theta)}{1-tan^2(\theta)}=tan(2\theta)\qquad thus\\ \quad \\ \quad \\ \cfrac{\frac{2tan(\theta)}{1-tan^2(\theta)}}{4}\implies \cfrac{2tan(\theta)}{1-tan^2(\theta)}\cdot \cfrac{1}{4}\implies \cfrac{tan(\theta)}{2[1-tan^2(\theta)]}\)